OpenStudy (anonymous):

The point P(3, 4) is on the terminal side of θ. Evaluate tan θ.

OpenStudy (amistre64):

im sure there is a definition for the tangent function ....

OpenStudy (amistre64):

do you recall the 3 basic trig functions and their definitions?

OpenStudy (anonymous):

Sin, Cos, Tan, I dont know the definitions though

OpenStudy (amistre64):

on a normal looking xy graph; the defintions equate to: \[sin=\frac{y}{\sqrt{x^2+y^2}}\] \[cos=\frac{x}{\sqrt{x^2+y^2}}\] \[tan=\frac{y}{x}\]

OpenStudy (anonymous):

Okay, sooo...

OpenStudy (amistre64):

so take the point (x,y) that they give you, and use those values in the appropriate definition.

OpenStudy (anonymous):

Example? I take this class online, you are basically my teacher right now

OpenStudy (amistre64):

spose they give you a point (a,b) and ask you to determine the value of say; tan you take the values presented in the point: x=a, y=b and then apply them to the appropriate definition; in this case tan = y/x, therefore; tan = b/a

OpenStudy (anonymous):

So tan= 3/4

OpenStudy (amistre64):

that was a very good attempt :) (x,y) = (3,4) y = 4 x = 3 tan = 4/3

OpenStudy (anonymous):

oh lol Sorry I didnt really read it I just assumed x/y sorry

OpenStudy (amistre64):

youll do fine :) good luck